OpenStudy (anonymous):

I need a solution step by step d/dx Derivative of trigonometric functions y=cot x - tan x

OpenStudy (fibonaccichick666):

so, do you happen to have the derivative of cotangent memorized?

OpenStudy (fibonaccichick666):

If not, check out this page(if we do not want to have to derive it every time, it is beneficial to have them memorized) http://www.sosmath.com/calculus/diff/der03/der03.html

OpenStudy (jdoe0001):

\(\bf \cfrac{d}{dx}[cot(\theta)]=-csc^2(\theta)\qquad \cfrac{d}{dx}[tan(\theta)]=sec^2(\theta)\)

OpenStudy (jdoe0001):

well.. or \(\bf \cfrac{d}{dx}[cot(x)]=-csc^2(x)\qquad \cfrac{d}{dx}[tan(x)]=sec^2(x)\) for that matter

OpenStudy (fibonaccichick666):

@pearlss can you please respond?

OpenStudy (jdoe0001):

or blink twice for yes or blnk once for no or blink three times for maybe